# Layered, Overlapping, and Inconsistent: A Large-Scale Analysis of the Multiple Privacy Policies and Controls of U.S. Banks

### v1

@@ 5.3.4. Third-Party Cookie Practices
Our cookie script analysis revealed that 1,454 banks’ websites (70.1%) contained third-party cookies, and 1,252 of them (60.4%) specifically included third-party marketing cookies (see Table 6). These findings sharply contrast with banks’ respective disclosures: only 184 websites (14.7% of 1,252 banks with third-party marketing cookies) mentioned any practice of allowing marketing cookies. Interestingly, we also found that of the 225 banks that disclosed allowing marketing third-party cookies, 41 (18.2%) did not actually have such third-party cookies on their website. This discrepancy may stem from vague or incomplete disclosures. Banks often fail to specify who receives cookie data, making it difficult to determine the true scope of their sharing practices or whether third parties are involved in marketing activities.
Table 7. CCPA disclosure and opt-out methods (n=2073).
@@ 5.3.5. CCPA Privacy Controls
We assessed whether the banks that are required to provide CCPA do-not-sell/share opt-out based on their sharing disclosures provided an opt-out link and responded to GPC signals (see Table 7). 200 banks explicitly stated that they do not sell/share personal information under CCPA. Interestingly, several of these still provided an opt-out link (18) and respected GPC signals (21). Either these banks are taking extra precautions to avoid being perceived as non-compliant with the CCPA, or their explicit assertion to not share/sell may not reflect their actual practices. Of the 45 banks that did acknowledge to sell/share data under CCPA, 22 (48.9%) failed to implement required opt-outs. Only 13 (28.9%) provided an opt-out link and 22 (48.9%) respected GPC signals. We also found a few banks that made no statements about selling/sharing under CCPA in their policies, yet still provided a do-not-sell/share opt-out link (10) or respected GPC signals (17). This inconsistency raises concerns that these banks may be engaging in CCPA-covered data sharing without transparently disclosing such practices in their privacy policies, or may be implementing opt-outs they are not required to provide.
RQ3 Summary. Although banks mostly implemented the required opt-outs for GLBA, half of the banks that disclosed sale/sharing of personal information under CCPA failed to implement the required opt-outs. Furthermore, while 1,252 banks (60.4%) allowed marketing third-party cookies on their websites, only 184 (14.7% of 1,252) disclosed this in their privacy policies.
@@ 6.1. Key Findings and Implications
Clarity of GLBA notice eroded by additional policies. In addition to the legally mandated GLBA notice, about 44% of the largest U.S. banks we examined provided at least one additional privacy policy, covering a wider range of data practices related to online and mobile services. Combining all privacy policies per bank, we also found that length and complexity increase with bank size. This proliferation of privacy statements, particularly the coexistence of GLBA notices and other privacy policies, reflects banks’ efforts to comply with an increasingly fragmented regulatory privacy environment. While the GLBA short-form notice was designed to enhance transparency and facilitate compliance, its narrow scope on financial personal information has not kept pace with the digital data practices banks now engage in. As a result, the GLBA short-form notice, though designed to be concise and user-friendly, may no longer serve as the primary or most informative privacy disclosure for consumers.
@@
Inconsistencies due to GLBA’s narrow scope. When a bank stated “no” for any GLBA sharing purpose, with a 15–27% likelihood, the same bank made related affirmative sharing statements elsewhere in its other privacy policies, such as allowing third-party marketing or analytics cookies. Overall, of 915 banks that provided a GLBA notice and at least one other policy, 55.2% banks’ policies contained at least one such inconsistency. Rather than legal noncompliance, this phenomenon is rooted in GLBA’s narrow scope on both financial personal information and particular sharing practices. Yet these legal nuances are likely difficult to recognize for consumers: the GLBA notice uses the same generic term “personal information” as other policies, and offers only five examples to illustrate what the personal information includes without clarifying what falls outside its scope (e.g., online behavioral tracking, location, etc.). As a result, consumers may reasonably interpret a “no” in the GLBA table as a denial of sharing, while other policies reveal otherwise, rendering the GLBA notice uninformative at best and misleading at worst.
@@
Inconsistencies due to CCPA’s protective impact. We observed a less concerning but still notable pattern in the opposite direction: when a bank indicated ”yes” to any GLBA sharing purpose, with a 20-37% likelihood, the same bank disclosed not sharing with third parties in its other privacy policies. Many banks explicitly limited such sharing for California residents, which highlights the CCPA’s effectiveness in limiting data exposure for consumers. Other laws, including GLBA, should also provide stronger protections against third-party sharing beyond requiring transparency.
@@
Limited and unclear privacy opt-outs. We found a substantial gap between banks’ disclosed third-party sharing and the opt-outs they provided, particularly in relation to CCPA opt-outs and cookie practices. Among the websites that disclosed allowing third-party marketing or analytics cookies, only 13.3% implemented cookie controls. These controls also varied widely, with over 40 different cookie labels used. Furthermore, many banks’ websites used third-party marketing cookies without disclosing it. Concerningly, half of the banks that disclosed the sale or sharing of personal information under CCPA failed to provide the required opt-out link or honor GPC signals. These findings demonstrate a lack of regulation and enforcement regarding third-party tracking on banks’ websites.
@@ 6.2. Public Policy Recommendations
Expand the data-sharing table. Furthermore, the GLBA data-sharing table should be updated to reflect today’s data practices, especially online tracking, and broader categories of data sharing. The GLBA model notice could integrate a list of commonly shared personal information types that consumers care about most based on the extensive privacy research conducted since the GLBA notice was developed (e.g., behavioral, location).
@@
Standardize and reconcile privacy policies across laws. The multiple privacy policies provided by the same bank underscore the urgent need to modernize the layered privacy regulation framework to match current practices. Regulators and industry should work toward standardization to reduce redundancy and inconsistency across privacy policies. This includes unifying terminology (e.g., replacing outdated terms like GLBA’s “non-public information” with clearer and consistent definitions of “personal information”) and harmonizing usable opt-out designs (e.g., modeled after the “Your Privacy Choices” CCPA opt-out link).
@@ 7. Conclusion
Our analysis of the 2,073 largest U.S. bank websites revealed a fragmented privacy disclosure environment for consumers. Nearly half of the banks provided multiple privacy policies, which are long, difficult to read, and especially complex for larger banks. We identified two key inconsistencies in third-party marketing-related disclosures. The concerning cases are where a bank stated “no” to sharing in GLBA notice but disclosed sharing elsewhere, which may mislead consumers. We also found limited alignment with CCPA-required opt-out controls and inadequate support for third-party cookie opt-outs. These findings highlight the difficulties that the layered and overlapping regulatory requirements bring to both banks and consumers. To restore transparency, we suggest that privacy disclosure requirements and opt-out controls must be harmonized across laws.
@@ Acknowledgements.
This research has been partially supported by the National Science Foundation under Award No. 2105734. We thank Lee Matheson, Shomir Wilson, Shahriar Shayesteh, Mukund Srinath, Maaz Bin Musa, Aysun Öğüt, and Eera Bhatt for their help and feedback and Nick Feamster for supporting Van Tran’s participation in the project. The authors used Grammarly for proofreading.

### v2

@@ 5.3.4. Third-party cookie practices
Our cookie script analysis revealed that 1,454 banks’ websites (70.1%) contained third-party cookies, and 1,252 of them (60.4%) specifically included third-party marketing cookies (see Table 6). However, among the 1,252 banks with third-party marketing cookies, only 184 websites (14.7%) disclosed any practice of allowing marketing cookies. Interestingly, we also found that of the 225 banks that disclosed allowing marketing third-party cookies, 41 (18.2%) did not actually have such third-party cookies on their website. This discrepancy may stem from vague or incomplete disclosures. Banks often fail to specify who receives cookie data, making it difficult for consumers to determine the true scope of their sharing practices or whether third parties are involved in marketing activities.
Table 7. CCPA disclosure and opt-out methods (n=2,073).
@@ 5.3.5. CCPA privacy controls
We assessed whether the banks that are required to provide CCPA do-not-sell/share opt-out based on their sharing disclosures provided an opt-out link and responded to GPC signals (see Table 7). 200 banks explicitly stated that they do not sell/share personal information under CCPA. Interestingly, several of these still provided an opt-out link (18) and respected GPC signals (21). Either these banks are taking extra precautions to avoid being perceived as non-compliant with the CCPA, or their explicit assertion not to share/sell may not reflect their actual practices. Of the 45 banks that did acknowledge to sell/share data under CCPA, 22 (48.9%) failed to implement required opt-outs. Only 13 (28.9%) provided an opt-out link, and 22 (48.9%) respected GPC signals. We also found a few banks that made no statements about selling/sharing under CCPA in their policies, yet still provided a do-not-sell/share opt-out link (10) or respected GPC signals (17). This inconsistency raises concerns that these banks may be engaging in CCPA-covered data sharing without transparently disclosing such practices in their privacy policies, or may be implementing opt-outs they are not required to provide.
RQ3 summary. Although banks mostly implemented the required opt-outs for GLBA, half of the banks that disclosed the sale/sharing of personal information under CCPA failed to implement the required opt-outs. Sixty percent of banks allowed marketing third-party cookies on their websites, and yet only fifteen percent of them disclosed this in their privacy policies. We also found that the naming of cookie controls and the labeling of different cookie types were both highly varied, which may confuse consumers and prevent them from effectively opting out.
@@ 6.1. Key Findings and Implications
Clarity of GLBA notice eroded by additional policies. Compared to Cranor et al. (Cranor et al., 2016) a decade ago, we found that fewer banks disclose data sharing and more provide required opt-outs in their GLBA notice. Our analysis of a more complete set of privacy notices that reflect various current regulatory requirements and disclosure practices also showed that about 44% of the largest U.S. banks we examined provided at least one other privacy policy in addition to the legally mandated GLBA notice. With these additional privacy policies that cover a wider range of data practices related to online and mobile services, the length and complexity of a bank’s privacy disclosures increase with bank size. This proliferation of privacy statements, particularly the coexistence of GLBA notices and other privacy policies, reflects banks’ efforts to comply with an increasingly fragmented regulatory privacy environment. While the GLBA short-form notice was designed to enhance transparency and facilitate compliance, its narrow scope on financial personal information has not kept pace with the digital data practices banks now engage in. As a result, the GLBA short-form notice, though designed to be concise and user-friendly, may no longer serve as the primary or most informative privacy disclosure for consumers.
@@
Inconsistencies due to GLBA’s narrow scope. When a bank stated “no” for any GLBA sharing purpose, with a 15–27% likelihood, the same bank made related affirmative sharing statements elsewhere in its other privacy policies, such as allowing third-party marketing or analytics cookies, with marketing/advertising disclosures more common than analytics/research. Overall, of 915 banks that provided a GLBA notice and at least one other policy, 53.8% banks’ policies contained at least one such inconsistency. Rather than legal noncompliance, this phenomenon is rooted in GLBA’s narrow scope on both financial personal information and particular sharing practices. Yet these legal nuances are likely difficult to recognize for consumers: the GLBA notice uses the same generic term “personal information” as other policies, and offers only five examples to illustrate what the personal information includes without clarifying what falls outside its scope (e.g., online behavioral tracking, location, etc.). As a result, consumers may reasonably interpret a “no” in the GLBA table as a denial of sharing, while other policies reveal otherwise, rendering the GLBA notice uninformative at best and misleading at worst.
@@
Inconsistencies due to CCPA’s protective impact. We observed a less concerning but still notable pattern in the opposite direction: when a bank indicated “yes” to any GLBA sharing purpose, with a 20-37% likelihood, the same bank disclosed not sharing with third parties in its other privacy policies. Many banks explicitly limited such sharing for California residents, which highlights the CCPA’s effectiveness in limiting data exposure for California consumers at least. Other laws, including GLBA, should also provide stronger protections against third-party sharing beyond requiring transparency.
@@
Limited and unclear privacy opt-outs. We found a substantial gap between banks’ disclosed third-party sharing and the opt-outs they provided, particularly in relation to CCPA opt-outs and cookie practices. Among banks’ websites that disclosed allowing third-party marketing or analytics cookies, only 13.3% implemented cookie controls. These controls also varied widely, with over 40 different cookie labels used. Furthermore, many banks’ websites used third-party marketing cookies without disclosing it. Notably, half of the banks that disclosed the sale or sharing of personal information under CCPA failed to provide the required opt-out link or honor GPC signals. These findings demonstrate a lack of regulation and enforcement regarding third-party tracking on banks’ websites.
Implications for future research. The specific forms of inconsistency we identified across privacy policies under different regulations, together with the different types of sharing statements, may inform future user studies on how overlapping policies and controls affect user understanding as well as the efficacy of policies and controls. Our collection of documents and annotated segments, which we made available as artifacts, could support training classifiers to identify sharing-related disclosures across policy types beyond those covered by existing resources (e.g., OPP-15 (Wilson et al., 2016)). Such tools may additionally support longitudinal monitoring of privacy policy changes, compliance audits of privacy practices, and analyses in other sectors with a more automated approach.
@@ 6.2. Public Policy Recommendations
Expand the data-sharing table. Furthermore, the GLBA data-sharing table should be updated to reflect today’s data practices, especially online tracking, and broader categories of data sharing. The GLBA model notice could integrate a list of commonly shared personal information types that consumers care about most based on the extensive privacy research conducted since the GLBA notice was developed (e.g., behavioral tracking, location, etc.).
@@
Standardize and reconcile privacy policies across laws. The multiple privacy policies provided by the same bank underscores the urgent need to modernize the layered privacy regulation framework to match current practices. Regulators and industry should work toward standardization to reduce redundancy and inconsistency across privacy policies. This includes unifying terminology (e.g., replacing outdated terms like GLBA’s “non-public information” with clearer and consistent definitions of “personal information”) and harmonizing usable opt-out designs (e.g., modeled after the “Your Privacy Choices” CCPA opt-out link).
@@ 7. Conclusion
Our analysis of the 2,073 largest U.S. bank websites revealed a fragmented privacy disclosure environment for consumers. Nearly half of the banks provided multiple privacy policies, which are long, difficult to read, and especially complex for larger banks. We identified two key inconsistencies in third-party marketing-related disclosures. The concerning cases are where a bank stated “no” to sharing in the GLBA notice but disclosed sharing elsewhere, which may mislead consumers. We also found limited alignment with CCPA-required opt-out controls and inadequate support for third-party cookie opt-outs. These findings highlight the difficulties that the layered and overlapping regulatory requirements bring to both banks and consumers. To restore transparency, we suggest that privacy disclosure requirements and opt-out controls must be harmonized across laws.
@@ Acknowledgements.
This research has been partially supported by the National Science Foundation under Award No. 2105734; Van Tran has additionally been supported by the National Science Foundation under Award No. 2334996. We thank Shomir Wilson, Shahriar Shayesteh, Kirsten Martin, Mukund Srinath, Maaz Bin Musa, Aysun Öğüt, Eera Bhatt, Lee Matheson, and Sumit Asthana for their input and feedback. The authors used Grammarly for proofreading.